ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

Film

1991 MS Classification: 01, 00 Näätänen, Marjatta, Director, University of Helsinki, Finland Women and Mathematics Across Cultures This film, produced in 1996, briefly introduces the organization European Women in Mathematics (EWM), provides some statistics, and allows four women mathematicians to share their personal experiences about the impact of cultural differences on the status of women in the profession. Its making was motivated by questions such as: Why are there many more women mathematicians in Italy and Brazil than in Norway, England or Germany? What is it about Latin culture that encourages women mathematicians, while that of Northern Europe mitigates against them? The film was directed by Näätänen in collaboration with Bodil Branner (Technical University of Denmark), Kari Hag (UNIT - NTH, Trondheim, Norway), and Caroline Series (University of Warwick, UK).